The ages of Ramesh and Rahim are in the ratio 5:7.If Ramesh were 9years older and Rahim 9years younger,the age of Ramesh would have been twice the age of Rahim find there ages.
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

The ages of Ramesh and Rahim are in the ratio 5:7.If Ramesh were 9years older and Rahim 9years younger,the age of Ramesh would have been twice the age of Rahim.
-----------------
Equations:
m/h = 5/7
(m+9) = 2(h-9)
----
Rearrange:
7m - 5h = 0
m - 2h = -27
---------------
Multiply thru 2nd equation by 7:
-------
7m - 5h = 0
7m - 14h = -189
---------------------
Subtract and solve for "h";
9h = 189
h = 21 (Rahim's age now)
---
Solve for "m":
m - 2h = -27
m - 2*21 = -27
m = -27+42
m = 15 (Ramesh's age now)
=========================
Cheers,
Stan H.
===========

Answer by lwsshak3(11628)

The ages of Ramesh and Rahim are in the ratio 5:7.If Ramesh were 9years older and Rahim 9years younger,the age of Ramesh would have been twice the age of Rahim find there ages.
..
let x=Age of Ramish
let y=Age of Rahim
x/y=5/7
5y=7x
y=7x/5
..
Ramish 9 years older=x+9
Rahim 9 years younger=y-9
(x+9)/(y-9)=2
x+9=2y-18
substitute y=7x/5
x+9=2(7x/5)-18
5x+45=14x-90
-9x=-135
x=15
y=7*15/5=21
ans:
Age of Ramish=15
Age of Rahim =21